# A Study on the Calibration of In-context Learning

### 11 Dec 2023 v2

@@ 4.3 Qualitative Results
Figure 6: The comparison of calibration errors before and after applying different recalibration strategies.
@@
We showcase that in SST-2 and Open Book QA (4-shot), both ACC and ECE of LLaMA increase as the model size increases (Fig. 2). Moreover, it’s observed that confidence scores tend to concentrate on high values as we enlarge the model size. In particular, the confidence of almost all predictions are above 0.8 for the 30B model.
@@ 4.4 Ablation Studies
For case studies, we research how miscalibration can impact the selective classification of LLMs, where models are supposed to abstain from uncertain predictions in high-stakes settings.
@@ 5 Discussion and Concluding Remarks
In our investigation of the token-level calibration of in-context learning in contemporary Large Language Models (LLMs), we delineate the intricate balance between ICL performance and calibration. Our findings underscore the importance of being circumspect in model deployment, as maximizing ICL performance does not invariably translate to improved calibration. As LMs continue to evolve and gain more capabilities such as having long enough context windows that can include the whole training set as in-context examples for some downstream tasks, our result can be pedagogical when users would like to examine their uncertainty through prediction probabilities. Moreover, the work suggests the following future directions:
@@
In this work, we observe that existing scaling recalibration methods cannot fully resolve the miscalibration issues of ICL, so better understanding and mitigation strategies are needed. A potential approach can be leveraging transparency tools and studying whether predictable errors exist during text generation.
@@
Our findings indicate that in multi-choice or multi-class classification tasks, even though the calibration of answer tokens may deteriorate in high-performance settings, there may be a positive correlation between accuracy and the correctness of explanations in reasoning tasks. This suggests potential avenues for future research such as exploring strategies such as the use of hedging words to express uncertainty and examining their relationship with predictive performance.
@@ Acknowledgment
We thank Yu Bai, David Childers, Jean-Stanislas Denain for valuable feedback. Kempner Institute computing resources enabled this work.

### 02 Jan 2024 v3

@@ 4.3 Qualitative Results
Recall that we found significant miscalibration for reasoning with CoT settings, therefore we closely examine the poorly calibrated reasoning cases using the above plots (Fig. 2 and Fig. 6). Our results on 4-shot settings show that for the reasoning problems (Strategy QA, Commonsense QA, OpenBook QA, World Tree) we consider, models are consistently over-confident with ECEs above 0.15. Larger models are better both in both ACC and ECE but for OpenBook QA, calibration worsens as the model size increases. Moreover, it’s observed that confidence scores tend to concentrate on high values as we enlarge the model size. Especially in Commonsense QA and OpenBook QA, the confidence level of nearly all predictions of 13B and 30B models predominantly exceeds 0.8.
@@ 4.4 Ablation Studies
For case studies, we research how miscalibration can impact the selective classification of LMs, where models are supposed to abstain from uncertain predictions in high-stakes settings.
@@ 5 Discussion and Concluding Remarks
In our investigation of the token-level calibration of in-context learning in contemporary language models, we illustrate the intricate trade-off between ICL performance and calibration. Our findings underscore the importance of being circumspect in model deployment, as maximizing ICL performance does not invariably translate to improved calibration for low-shot and reasoning settings. As LMs continue to evolve and gain more capabilities such as having long enough context windows that can include the whole training set as in-context examples for some downstream tasks, our result can be pedagogical when users would like to examine their uncertainty through prediction probabilities. Moreover, the work suggests the following future directions:
@@
We observe that existing scaling recalibration methods cannot fully resolve the miscalibration issues of ICL, so better understanding and mitigation strategies are needed. A potential approach can be leveraging transparency tools and studying whether predictable errors exist during text generation.
@@
Our findings indicate that in multi-choice or multi-class classification tasks, even though the calibration of answer tokens may deteriorate in high-performance settings, there may be a positive correlation between accuracy and the correctness of explanations in reasoning tasks. This suggests potential avenues for future research in exploring strategies such as the use of hedging words to express uncertainty and examining their relationship with predictive performance.
Implications in assessing beliefs of LMs.
Previous works show that the expected calibration error would decrease monotonically as the number of ICL examples increases (Kadavath et al., 2022) when querying LMs for answer probabilities. However, we find that zero-shot performance might be weak for models less than 30B, and in low-shot settings, calibration errors can sometimes be even worse than zero-shot. This implies that a close examination and careful control of epistemic uncertainty and aleatoric uncertainty can be needed before deriving conclusions in truthfulness (Liu et al., 2023; Azaria and Mitchell, 2023) for low-shot settings.
@@ Acknowledgment
We thank Yu Bai, David Childers, Jean-Stanislas Denain for their valuable feedback. Kempner Institute computing resources enabled this work.
Appendix A Additional Experimental Details
We provide prompts we adopt for experiments in Tab.6. Additional reliability plots are shown in Fig. 6.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 6: Reliability plots and confidence histograms of LLaMA models on 4-shot reasoning tasks. Results of different sizes 7B (left), 13B (middle), and 30B (right) are plotted.
Data: $\mathcal{P}_{\theta}(\textbf{w})$: Original output of the classification model, $\mathcal{D}$: Training dataset, $\tau$: Temperature parameter, $k$: we use $k$-shot experimental settings, where during test the ICL prompts will consist of $k$ (sample, label) pairs.
Result: Adjusted probabilities after temperature scaling;
// Training process
// 00-shot: $(\mathbf{w}_{i},y_{i})$ is every training samples and corresponding label.
// $k$-shot: $\mathbf{w}_{i}=\{x_{1},y_{1},...,x_{k},y_{k},x_{i}\}$ uses $k$ prompt pairs.
// Fix $w$: the prompt in $\mathbf{w}_{i}=\{x_{1},y_{1},...,x_{k},y_{k},x_{i}\}$ will be used for all training instances and used during inference.
for each training sample $(\mathbf{w}_{i},y_{i})\in\mathcal{D}$ do
       Compute the original output: $z_{i}=\mathcal{P}_{\theta}(\mathbf{w}_{i};\theta)$;
       Compute the cross-entropy loss: $L_{i}=\text{CrossEntropy}(z_{i},y_{i})$;
end for
Compute the gradient of the loss to the temperature parameter: $\nabla_{\tau}\mathcal{L}=\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\nabla_{\tau}L_{i}$;
Update the temperature parameter using gradient descent: $\tau\leftarrow\tau-\eta\nabla_{\tau}\mathcal{L}$;
// Test time
for each test sample $\mathbf{x}_{j}$ do
       Compute the original output with prompt: $z_{j}=\mathcal{P}_{\theta}(\mathbf{x}_{j};\theta)$;
       Compute the adjusted output: $\hat{z}_{j}=\frac{z_{j}}{\tau}$;
       Compute the softmax probabilities: $\hat{p}_{j}=\text{Softmax}(\hat{z}_{j})$;
end for
Algorithm 1 Pseudocode for temperature scaling
Dataset Prompt Label
SST-2
Review: it may not be a great piece of filmmaking, but its power comes from its soul’s - eye view of how well-meaning patronizing masked a social injustice, at least as represented by this case .
Sentiment: Positive
Review: smith’s point is simple and obvious – people’s homes are extensions of themselves, and particularly eccentric people have particularly eccentric living spaces – but his subjects are charmers .
Sentiment:
Negative, Positive
CB
A: No, not really. I spend a lot of time with our income tax, though. especially, this year and last year. Um, I have been married for just a few years, so I’ve had to really switch around from the EZ form to the, uh, B: Schedule A. A: Right. B: Well, yeah. A: All the deductions and all that. B: Did you notice that when they passed the new simplified tax act, it seemed like it made everything harder?
question: when they passed the new simplified tax act it seemed like it made everything harder. true, false, or neither?
answer: true
There was a group of curious onlookers… Marie felt her legs give way beneath her, she sat down on the edge of the pavement, feet in the gutter, doubled-up, sick and winded as if someone had punched her in the stomach. She lifted up her head and looked again. She had watched scenes like this so often in detective films and police series on television that she could hardly believe that this was real life.
question: this was real life. true, false, or neither?
answer:
True, False, Neither
RTE
The main institutionalised forms of recognition for those who have made a significant contribution in the fields of physics, chemistry, medicine, literature, as well as for those working for peace (and more recently in the area of economics), are the Nobel prizes.
question: Nobel Peace Prize candidates have been chosen. True or False?
answer: False
Egypt on Thursday strongly criticized Israeli new Foreign Minister Avigdor Lieberman for his remarks that he refused to recognize the peace efforts initiated in 2007 in the U.S. city of Annapolis to restore the peace talks with the Palestinians, reported the state MENA news agency. Liebermanś remarks is "regrettable," Egyptian Foreign Ministry spokesman Hossam Zaki was quoted as saying, adding "his remarks are the first blow to the peace efforts to come from the Israeli new government."
question: Hossam Zaki is the new Foreign Minister of Israel. True or False?
answer:
True, False
Strategy QA
Question: Can spiders help eggplant farmers control parasites? Choose the answer from True and False.
Answer: The potato tuber moth is a parasite that targets the plant family Solanaceae, including eggplant Selenops radiatus is a spider genus in South Africa that effectively controls the potato tuber moth So, the answer is: True
Question: Is the voice of Genie from Disney’s Aladdin still alive? Choose the answer from True and False.
Answer:
True, False
Commonsense QA
"Question: Dan was a farmer with just one heifer. But that was okay, he only kept her for milk, and he didn’t think he’d find good farmland in a place as cold as where?
A arizona
B farm yard
C michigan
D german field
E dairy farm
Answer: Michigan is a state in the us where it precipitates throughout the year and areas, where it precipitates throughout the year, are generally cold. So the farmer thought he’d not find a good farmland in a place as cold as michigan. Enslaving heifers or other animals for their milk is wrong as they want to live free. All the places in the other options may not be cold. So, the answer is: C
Question: From where does a snowflake form?
A cloud
B snow storm
C billow
D air
E snowstorm
Answer:"
A, B, C, D, E
Table 6: Prompts used for text classification and reasoning tasks, with a single training example showcased per task for illustrative purposes. The right column displays corresponding labels. The prompting formats and labels for WorldTree and OpenBookQA are the same as those of the CommonsenseQA dataset.